zarkam21 (zarkam21):

Which of the following could be the ratio of the length of the longer leg of a 30-60-90 triangle to the length of its hypotenuse? Check all that apply. A. 2 √3 : 4 B. √3 : √3 C. √3 : 2 D. 1 : √3 E. 2: 2 √2 F. √2: √3

OpenStudy (denonakavro):

The 3 values in ascending order are: 1,sqrt3,2 shorter leg would be 1 longer leg would be sqrt 3 hypotenuse would be 2 So longleg/hypot = sqrt3/2 or sqrt3:2

zarkam21 (zarkam21):

So choice C and F?

zarkam21 (zarkam21):

So just c?

OpenStudy (denonakavro):

Only C

OpenStudy (phi):

notice that in a ratio, if you multiply both the top and bottom , we get a similar ratio for example \[ \frac{\sqrt{3}}{2} = \frac{2\sqrt{3}}{4}\] in other words, choice A is also a possible ratio of long side to hypotenuse in a 30-60-90 triangle

OpenStudy (radar):

A good point phi, the choices of both A and C were presented, wonder how the instructor would have treated an answer of A ??

OpenStudy (denonakavro):

radar I believe that would also be correct for the instructor, because it says check all that apply.